# Gravitational Lensing of Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter

### v3

@@ 1 Introduction
The equation of state for dark matter in galaxies (around the supermassive black hole in the galaxy center) is still a mystery. Recently, a perfect fluid dark matter (PFDM) model has attracted great interest [28, 29, 30]. Despite the simple form of PFDM model, it can reflect some non-trivial effects and properties of dark matter. In recent years, a number of studies investigate the PFDM effects on many aspects of black hole physics, and effective spacetime metrics for several kinds of black holes (such as Schwarzschild, RN, Kerr black holes) surrounded by PFDM have been successfully constructed [31, 32, 33, 34, 35]. Based on these effective spacetime metrics, the gravitational lensing, black hole shadow images, quasi-normal mode, thermodynamic properties of black hole system in PFDM medium can be extensively analyzed [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54].
@@ 2 Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter
where $F=\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$ is the conventional electromagnetic invariant in Maxwell theory and $G=\frac{1}{4}F_{\mu\nu}*F^{\mu\nu}$ (the $*F_{\mu\nu}$ is the dual form of the electromagnetic strength tensor $F_{\mu\nu}$). The above Lagrangian leads to the gravitational field equation
@@
$(T_{\text{EM}})_{\nu}^{\mu}$ $=$ $\frac{1}{4\pi}\cdot\text{diag}\bigg{\{}-\frac{Q^{2}}{2r^{4}}+\frac{aQ^{4}}{8r^{8}},\ -\frac{Q^{2}}{2r^{4}}+\frac{aQ^{4}}{8r^{8}},\ \frac{Q^{2}}{2r^{4}}-\frac{3aQ^{4}}{8r^{8}},\ \frac{Q^{2}}{2r^{4}}-\frac{3aQ^{4}}{8r^{8}}\bigg{\}},\ \ \text{Electromagnetic Field}$
@@
The motivation and detailed derivation of the metric component function $f(r)$ are reviewed in Appendix A. In the absence of dark matter (namely $\lambda_{\text{DM}}=0$), this metric represents an Euler-Heisenberg black hole with nonlinear electrodynamics effects. On the other hand, when $a=0$, the nonlinear electrodynamics effects vanish and the spacetime gives a charged Reissner-Nordström (RN) black hole surrounded by PFDM. In the cases of $\lambda_{\text{DM}}=0$ and $a=0$, the spacetime recovers to a simple RN black hole spacetime.
In the presence of PFDM and nonlinear electrodynamics effects, the horizon of charged black hole can be greatly influenced. It is commonly known that the classical charged black hole spacetime (such as RN spacetime) could produce naked singularities when electric charge exceeds its maximal allowed value. However, the Euler-Heisenberg black hole with a nonzero nonlinear electrodynamics parameter $a>0$ and a non-vanishing PFDM parameter $\lambda_{\text{DM}}\neq 0$ would no longer produce naked singularities for arbitrary black hole electric charge (see Appendix B). Therefore, in the rest of this work, it is safely to change black hole parameters in the calculation of gravitational lensing observables for Euler-Heisenberg black hole surrounded by PFDM, without worrying about the annoying naked singularities. The horizons of Euler-Heisenberg black hole surrounded by PFDM are discussed in Appendix B.
@@ 3.2 Gravitational Deflection Angle of Light
$\frac{d\phi}{dr}=-\frac{1}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}}}<0,\ \ \ \ \text{from source position $r=r_{\text{S}}$ to tuning point $r=r_{0}$}$
@@
[…]
@@ 3.3 Time Delay of Light
$\frac{dt}{dr}=-\frac{1}{f(r)\sqrt{1-b^{2}\cdot\frac{f(r)}{r^{2}}}}<0,\ \ \ \ \text{from source position $r=r_{\text{S}}$ to tuning point $r=r_{0}$}$
@@ 3.4 Massive Object’s Orbits and Precession Angle
[…]
@@ 3.5 Black Hole Shadow
$\mathcal{K}(r=r_{\text{unstable}})<0$$\Leftrightarrow$ $\ \ \frac{d^{2}V_{\text{eff}}(r)}{dr^{2}}\bigg{|}_{r=r_{\text{unstable}}}<0,$ (29a)
$\mathcal{K}(r=r_{\text{stable}})>0$$\Leftrightarrow$ $\ \ \frac{d^{2}V_{\text{eff}}(r)}{dr^{2}}\bigg{|}_{r=r_{\text{stable}}}>0.$ (29b)
@@ 4.1 Gravitational Deflection Angle of Light
Figure 3: The gravitational deflection angle of light for Euler-Heisenberg black hole surrounded by PFDM. The gravitational deflection angles are calculated numerically from expression (16) for the finite distance light source and observer. This figure highlights the influences of black hole electric charge, nonlinear electrodynamics effects and dark matter effects on gravitational deflection angles. (a) The upper left panel exhibits the influences from black hole electric charge $Q$, with the nonlinear electrodynamics parameter $a=0.5M^{2}$ and PFDM parameter $\lambda_{\text{DM}}=0.2M$. (b) The upper right panel shows the influences from nonlinear electrodynamics effects / QED effects, with the black hole charge $Q=0.5M$ and PFDM parameter $\lambda_{\text{DM}}=0.2M$. (c) The lower left panel highlights the influences from dark matter in the absence of electric charge and nonlinear electrodynamics effects (with $Q=0$ and $a=0$). (d) The lower right panel highlights the influences from dark matter in the presence of nonlinear electrodynamics effects, with the black hole charge $Q=0.5M$ and nonlinear electrodynamics parameter $a=0.5M^{2}$. In all four panels of this figure, the horizontal axis labels the impact parameter in photon orbit, which is increased from $b=10M$ to $b=100M$. The positions of light source and observer are located at a certain distance $r_{\text{O}}=r_{\text{S}}=200M$.
@@
The gravitational deflection angles of light for Euler-Heisenberg black hole surrounded by PFDM are given in figure 3. They are calculated numerically from expression (16) for light source and observer located at a finite distance region. In figure 3, the upper left and upper right panels highlight the influences from black hole electric charge $Q$ and the nonlinear electrodynamics effects / QED effects (parameterized by $a$), while the lower panels highlight the influences from dark matter parameterized by $\lambda_{\text{DM}}$. To show the gravitational deflection in strong gravitational field cases, the minimal value of impact parameter is chosen as $b\sim 10M$, which indicates the photon orbit is very close to the black hole critical shadow radius.
@@
The upper left panel of figure 3 suggests that, for a constant PFDM parameter and nonlinear electrodynamics parameter, the gravitational deflection angles are slightly reduced when black hole’s electric charge becomes larger. But the effect coming from electric charge is small such that we must add a magnified subplot into upper left panel to separate the various curves corresponding to different black hole charges. The upper right panel of figure 3 suggests that the nonlinear electrodynamics parameter has no obvious influences on the gravitational deflection angles, even if a magnified subplot is employed. Comparing the upper and lower panels of figure 3, we can observe that both electric charge and nonlinear electrodynamics have minor impacts on gravitational deflection angle, while the most significant influences coming from the PFDM. From the lower panels of figure 3, it can be clearly manifested that the variation of PFDM parameter could dramatically change the gravitational deflection angle of light for charged Euler-Heisenberg black hole. Notably, the gravitational deflection angle reduces rapidly as the increasing of PFDM parameter $\lambda_{\text{DM}}$. This can be explained from the spacetime metric in expression (6) for Euler-Heisenberg black hole surrounded by PFDM. The dark matter effects contribute to the spacetime metric function $f(r)$ via the $-\lambda_{\text{DM}}/r$ term (if we naively omit the slowly varying logarithmic term $\ln\frac{r}{|\lambda_{\text{DM}}|}$ for simplicity), which is similar to the black hole’s mass term $2M/r$. This indicates that the PFDM parameter could play a role of “effective mass” in the spacetime metric  [^1] Strictly specking, because of the presence of logarithmic term, this “effective mass” should be $m_{\text{eff}}(r)=\lambda_{\text{DM}}\ln\frac{r}{|\lambda_{\text{DM}}|}$, rather than $m_{\text{eff}}=\lambda_{\text{DM}}$.. A positive PFDM parameter ($\lambda_{\text{DM}}>0$) contributes to a negative effective mass in the spacetime metric, which explains the rapid drop of the gravitational deflection angle when PFDM parameter becomes larger. Furthermore, combining the four panels of figure 3, it is also interesting to find that the gravitational deflection angles of light for Euler-Heisenberg black hole surrounded by PFDM (with nonzero electric charge $Q\neq 0$ and dark matter parameter $\lambda_{\text{DM}}\neq 0$) are smaller than those in Schwarzschild black hole cases (which correspond to the black solid line in the lower left panel of figure 3).
@@ 4.2 Time Delay of Light
The time delay results for Euler-Heisenberg black hole surrounded by PFDM are presented in figure 4. This figure plots the time delay of light measured by the conventional static time coordinate $t$. They are obtained numerically from expression (19) for light source and observer located at a finite distance region. In this figure, the time delays of light varied with respect to light source position $r_{\text{S}}$ are plotted. A similar plot can be carried out for time delays varied with respect to observer position $r_{\text{O}}$, while these results exhibit very similar behaviors with those in figure 4. So the plot of time delays varied with respect to observer position is not given in this work. In figure 4, the upper left and upper right panels highlight the influences from black hole electric charge $Q$ and the nonlinear electrodynamics effects (or QED effects) parameterized by $a$, while the lower panels highlight the influences from dark matter parameterized by $\lambda_{\text{DM}}$. In all four panels, the time delays of light all grow with the increasing of light source radius $r_{\text{S}}$, confirming that the time delay of light $\Delta T$ calculated using expression (19) is indeed a monotonically increasing function of light source radius $r_{\text{S}}$ (or observer radius $r_{\text{O}}$).
Figure 4: The gravitational time delay of light for Euler-Heisenberg black hole surrounded by PFDM. The time delays of light (which are measured by the conventional static time coordinate $t$) are calculated numerically from expression (19) for the finite distance light source and observer. This figure plots the time delay results, highlighting the influences coming from various black hole parameters. (a) The upper left panel shows the influences from black hole electric charge $Q$, with the nonlinear electrodynamics parameter $a=0.5M^{2}$ and dark matter parameter $\lambda_{\text{DM}}=0.2M$. (b) The upper right panel exhibits the influences from nonlinear electrodynamics effects / QED effects, with the black hole charge $Q=0.5M$ and PFDM parameter $\lambda_{\text{DM}}=0.2M$. (c) The lower left panel highlights the influences from dark matter in the absence of black hole charge and nonlinear electrodynamics effects (with $Q=0$ and $a=0$). (d) The lower right panel highlights the influences from dark matter in the presence of nonlinear electrodynamics effects, with the black hole charge $Q=0.5M$ and nonlinear electrodynamics parameter $a=0.5M^{2}$. In all four panels of this figure, the horizontal axis labels the position of light source varied from $r_{\text{S}}=100M$ to $r_{\text{S}}=1000M$, and the vertical axis labels the time delay of light $\Delta T$ in unit of black hole mass. The position of observer is chosen to be $r_{\text{O}}=200M$, and the impact parameter in photon orbit satisfies $b=10M$.
@@ 4.3 Massive Object’s Bound Orbits and Precession Angle
Figure 5: The precession angles and trajectories of massive particle’s bound orbits moving around the Euler-Heisenberg black hole in the presence of PFDM. The precession angles $\Delta\phi_{\text{precession}}$ plotted in this figure are calculated numerically from expression (24) for a given conserved angular momentum value, and the trajectories of orbits are obtained by numerically solving the differential equation (23). The left panels shows the variation of precession angle as the changes of total energy parameter $\varepsilon_{\text{energy}}=\frac{E^{2}-1}{2}$, while the right panels illustrate the trajectories of particle orbits (It should be noted that we have substituted $1/2$ in the both side of equation (10) such that the energy parameter vanished at infinity, namely $\lim_{r\to\infty}\varepsilon_{\text{energy}}=0$). (a) The upper left panel plots the precession angle of orbits affects by black hole charge $Q$. (b) The upper right panel illustrates the trajectories of orbits for different black hole charge values when total energy equals to $\varepsilon_{\text{energy}}=-0.01$ (c) The middle left panel shows the precession angle of orbits affects by dark mater (parameterized by $\lambda_{\text{DM}}$). (d) The middle right panel illustrates the trajectories of orbits for different PFDM parameter values with total energy to be $\varepsilon_{\text{energy}}=-0.01$ (however, in the case of $\lambda_{\text{DM}}=0.2M$, the bound orbit does not exist because the local minimum of effective potential $V_{\text{min}}<\varepsilon_{\text{energy}}=-0.01$, see table 1) (e) The lower left panel shows the precession angle of orbits influenced by nonlinear electrodynamics effects / QED effects (parameterized by $a$). (f) The lower right panel illustrates the trajectories of orbits for different nonlinear electrodynamics parameter values with total energy to be $\varepsilon_{\text{energy}}=-0.01$
The figure 5 displays numerical results on massive object’s bound orbits around the Euler-Heisenberg black hole surround by PFDM, including the precession angles in the left panels and the trajectories of orbits in the right panels. The trajectories of massive object’s orbits are obtained by numerically solving the ordinary differential equation (23), and the precession angles of orbits are calculated from the integral in expression (24). The upper, middle and lower panels of figure 5 highlight the influences from black hole electric charge, dark matter (parameterized by PFDM parameter $\lambda_{\text{DM}}$), and nonlinear electrodynamics effects / QED effects (parameterized by nonlinear electrodynamics parameter $a$) correspondingly. In this figure, we present results on the trajectory of orbits and their precession angles for a selected angular momentum value $J=4M$. The variation trend of precession angles corresponding to other angular momentum values is similar to that for $J=4M$ presented in figure 5  [^2] For other angular momentum values, the proper region for conserved energy where bound orbits of massive objects could exist is different from those plotted in figure 5..

### v4

@@ 1 Introduction
The equation of state for dark matter in galaxies (around the supermassive black hole in the galaxy center) is still a mystery. Recently, a perfect fluid dark matter (PFDM) model has attracted great interest [28, 29, 30]. Despite the simple form of PFDM model, it can reflect some non-trivial effects and properties of dark matter. In recent years, a number of studies investigate the PFDM effects on many aspects of black hole physics, and effective spacetime metrics for several kinds of black holes (such as Schwarzschild, Reissner-Nordström, Kerr black holes) surrounded by PFDM have been successfully constructed [31, 32, 33, 34, 35]. Based on these effective spacetime metrics, the gravitational lensing, black hole shadow images, quasi-normal mode, thermodynamic properties of black hole system in PFDM medium can be extensively analyzed [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54].
@@ 2 Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter
where $F=\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$ is the conventional electromagnetic invariant in Maxwell theory and $G=\frac{1}{4}F_{\mu\nu}*F^{\mu\nu}$ (the $F_{\mu\nu}$ is the dual form of the electromagnetic strength tensor $F_{\mu\nu}$). The above Lagrangian leads to the gravitational field equation
@@
$(T_{\text{EM}})_{\nu}^{\mu}$ $=$ $\frac{1}{4\pi}\cdot\text{diag}\bigg\{-\frac{Q^{2}}{2r^{4}}+\frac{aQ^{4}}{8r^{8}},\ -\frac{Q^{2}}{2r^{4}}+\frac{aQ^{4}}{8r^{8}},\ \frac{Q^{2}}{2r^{4}}-\frac{3aQ^{4}}{8r^{8}},\ \frac{Q^{2}}{2r^{4}}-\frac{3aQ^{4}}{8r^{8}}\bigg\},\ \ \text{Electromagnetic Field}\ \ \ \$
@@
In the absence of dark matter (namely $\lambda_{\text{DM}}=0$), this metric represents an Euler-Heisenberg black hole with nonlinear electrodynamics effects. On the other hand, when $a=0$, the nonlinear electrodynamics effects vanish and the spacetime gives a charged Reissner-Nordström (RN) black hole surrounded by PFDM. In the cases of $\lambda_{\text{DM}}=0$ and $a=0$, the spacetime recovers to a simple RN black hole spacetime.
In the presence of PFDM and nonlinear electrodynamics effects, the horizon of charged black hole can be greatly influenced. It is commonly known that the classical charged black hole spacetime (such as RN spacetime) could produce naked singularities when electric charge exceeds its maximal allowed value. The Euler-Heisenberg black hole with a positive nonlinear electrodynamics parameter $a>0$ and a non-vanishing PFDM parameter $\lambda_{\text{DM}}\neq 0$ would no longer produce naked singularities for arbitrary black hole electric charge (see Appendix A). However, the Euler-Heisenberg black hole with a negative nonlinear electrodynamics parameter $a<0$ could still result in naked singularity. The horizons of Euler-Heisenberg black hole surrounded by PFDM are discussed in Appendix A.
@@ 3.2 Gravitational Deflection Angle of Light
$\frac{d\phi}{dr}=-\frac{1}{r^{2}\sqrt{\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}}}<0,\ \ \ \ \text{from source position $r=r_{\text{S}}$ to tuning point $r=r_{0}$}\ \ \ \$
@@
$=$ $\Psi_{\text{O}}-\Psi_{\text{S}}+\int_{r_{\text{S}}}^{r_{0}}\frac{d\phi}{dr}dr+\int_{r_{0}}^{r_{\text{O}}}\frac{d\phi}{dr}dr$
@@ 3.3 Time Delay of Light
$\frac{dt}{dr}=-\frac{1}{f(r)\sqrt{1-b^{2}\cdot\frac{f(r)}{r^{2}}}}<0,\ \ \ \ \text{from source position $r=r_{\text{S}}$ to tuning point $r=r_{0}$}\ \ \$
@@ 3.4 Massive Object’s Orbits and Precession Angle
$\Delta\phi_{\text{precession}}$ $=$ $\int_{r_{\text{min}}}^{r_{max}}\frac{d\phi}{dr}dr+\int_{r_{\text{max}}}^{r_{min}}\frac{d\phi}{dr}dr-2\pi$ (24)
@@ 3.5 Black Hole Shadow
$\mathcal{K}(r=r_{\text{unstable}})<0\ \ \displaystyle\Leftrightarrow$ $\ \ \frac{d^{2}V_{\text{eff}}(r)}{dr^{2}}\bigg|_{r=r_{\text{unstable}}}<0,$ (29a)
$\mathcal{K}(r=r_{\text{stable}})>0\ \ \displaystyle\Leftrightarrow$ $\ \ \frac{d^{2}V_{\text{eff}}(r)}{dr^{2}}\bigg|_{r=r_{\text{stable}}}>0.$ (29b)
@@ 4.1 Gravitational Deflection Angle of Light
Figure 3: The gravitational deflection angle of light for Euler-Heisenberg black hole surrounded by PFDM. (a) The upper left panel exhibits the influences from black hole electric charge $Q$. (b) The upper right panel shows the influences from nonlinear electrodynamics effects / QED effects. (c) The lower left panel highlights the influences from dark matter in the absence of electric charge and nonlinear electrodynamics effects. (d) The lower right panel highlights the influences from dark matter in the presence of nonlinear electrodynamics effects. In all panels, the horizontal axis labels the impact parameter in photon orbit. The positions of light source and observer are located at a certain distance $r_{\text{O}}=r_{\text{S}}=200M$.
@@
The gravitational deflection angles of light for Euler-Heisenberg black hole surrounded by PFDM are given in figure 3. They are calculated numerically from expression (16) for light source and observer located at a finite distance region. The multiple panels of figure 3 highlight the influences of black hole electric charge, nonlinear electrodynamics effects and dark matter effects on gravitational deflection angles. The upper left and upper right panels highlight the influences from black hole electric charge $Q$ and the nonlinear electrodynamics effects / QED effects (parameterized by $a$), while the lower panels highlight the influences from dark matter parameterized by $\lambda_{\text{DM}}$. To show the gravitational deflection in strong gravitational field cases, the minimal value of impact parameter is chosen as $b\sim 10M$, which indicates the photon orbit is very close to the black hole critical shadow radius.
@@
The upper left panel of figure 3 suggests that, for a constant PFDM parameter and nonlinear electrodynamics parameter, the gravitational deflection angles are slightly reduced when black hole’s electric charge becomes larger. But the effect coming from electric charge is small such that we must add a magnified subplot into upper left panel to separate the various curves corresponding to different black hole charges. The upper right panel of figure 3 suggests that the nonlinear electrodynamics parameter has no obvious influences on the gravitational deflection angles, even if a magnified subplot is employed. Comparing the upper and lower panels of figure 3, we can observe that both electric charge and nonlinear electrodynamics have minor impacts on gravitational deflection angle, while the most significant influences coming from the PFDM. From the lower panels of figure 3, it can be clearly manifested that the variation of PFDM parameter could dramatically change the gravitational deflection angle of light for charged Euler-Heisenberg black hole. Notably, the gravitational deflection angle reduces rapidly as the increasing of PFDM parameter $\lambda_{\text{DM}}$. This can be explained from the spacetime metric in expression (6) for Euler-Heisenberg black hole surrounded by PFDM. The dark matter effects contribute to the spacetime metric function $f(r)$ via the $\lambda_{\text{DM}}/r$ term (if we naively omit the slowly varying logarithmic term $\ln\frac{r}{|\lambda_{\text{DM}}|}$ for simplicity), which is similar to the black hole’s mass term $2M/r$. This indicates that the PFDM parameter could play a role of “effective mass” in the spacetime metric  [^1] Strictly specking, because of the presence of logarithmic term, this “effective mass” should be $m_{\text{eff}}(r)=\lambda_{\text{DM}}\ln\frac{r}{|\lambda_{\text{DM}}|}$, rather than $m_{\text{eff}}=\lambda_{\text{DM}}$.. A positive PFDM parameter ($\lambda_{\text{DM}}>0$) contributes to a negative effective mass in the spacetime metric, which explains the rapid drop of the gravitational deflection angle when PFDM parameter becomes larger. Furthermore, combining the four panels of figure 3, it is also interesting to find that the gravitational deflection angles of light for Euler-Heisenberg black hole surrounded by PFDM (with nonzero electric charge $Q\neq 0$ and dark matter parameter $\lambda_{\text{DM}}\neq 0$) are smaller than those in Schwarzschild black hole cases (which correspond to the black solid line in the lower left panel of figure 3).
@@ 4.2 Time Delay of Light
The time delay results for Euler-Heisenberg black hole surrounded by PFDM are presented in figure 4. This figure plots the time delay of light measured by the conventional static time coordinate $t$, which is calculated numerically from expression (19) for light source and observer located at a finite distance region. In this figure, the time delays of light varied with respect to light source position $r_{\text{S}}$ are plotted. A similar plot can be carried out for time delays varied with respect to observer position $r_{\text{O}}$, while these results exhibit very similar behaviors with those in figure 4. So the plot of time delays varied with respect to observer position is not given in this work. The multiple panels of figure 4 highlight the influences coming from various black hole parameters. The upper left and upper right panels highlight the influences from black hole electric charge $Q$ and the nonlinear electrodynamics effects (or QED effects) parameterized by $a$, while the lower panels highlight the influences from dark matter parameterized by $\lambda_{\text{DM}}$. In all panels, the time delays of light all grow with the increasing of light source radius $r_{\text{S}}$, confirming that the time delay of light $\Delta T$ calculated using expression (19) is indeed a monotonically increasing function of light source radius $r_{\text{S}}$ (or observer radius $r_{\text{O}}$).
Figure 4: The gravitational time delay of light for Euler-Heisenberg black hole surrounded by PFDM. (a) The upper left panel shows the influences from black hole electric charge $Q$. (b) The upper right panel exhibits the influences from nonlinear electrodynamics effects / QED effects. (c) The lower left panel highlights the influences from dark matter in the absence of black hole charge and nonlinear electrodynamics effects. (d) The lower right panel highlights the influences from dark matter in the presence of nonlinear electrodynamics effects. In all panels, the horizontal axis labels the position of light source, and the vertical axis labels the time delay of light $\Delta T$ measured in unit of black hole mass.
@@ 4.3 Massive Object’s Bound Orbits and Precession Angle
Figure 5: The precession angles and trajectories of massive particle’s bound orbits moving around the Euler-Heisenberg black hole in the presence of PFDM. (a) The upper left panel plots the precession angle of orbits affects by black hole charge. (b) The upper right panel illustrates the trajectories of orbits for different black hole charge values. (c) The middle left panel gives the precession angle of orbits affects by dark mater. (d) The middle right panel illustrates the trajectories of orbits for different PFDM parameter values. (e) The lower left panel shows the precession angle of orbits influenced by nonlinear electrodynamics effects / QED effects. (f) The lower right panel illustrates the trajectories of orbits for different nonlinear electrodynamics parameter values.
The figure 5 displays numerical results on massive object’s bound orbits around the Euler-Heisenberg black hole surround by PFDM, including the precession angles in the left panels and the trajectories of orbits in the right panels. The left panels of this figure present the variation of precession angle as the changes of total energy parameter $\varepsilon_{\text{energy}}=\frac{E^{2}-1}{2}$  [^2] It should be noted that we have substituted $1/2$ in the both side of equation (10) such that the energy parameter vanished at infinity, namely $\lim_{r\to\infty}\varepsilon_{\text{energy}}=0$., while the right panels illustrate the trajectories of particle orbits for a given energy parameter $\varepsilon_{\text{energy}}=-0.01$  [^3] In the middle right panel of figure 5, the bound orbit does not exist for the $\lambda_{\text{DM}}=0.2M$ case, because the local minimum of effective potential $V_{\text{min}}<\varepsilon_{\text{energy}}=-0.01$. So the red dotted curve, which represent the particle orbit for the $\lambda_{\text{DM}}=0.2M$ case, is disappear in the middle right panel. The local minimum of effective potential is given in table 1. The trajectories of massive object’s orbits are obtained by numerically solving the ordinary differential equation (23), and the precession angles of orbits are calculated from the integral in expression (24). The upper, middle and lower panels of figure 5 highlight the influences from black hole electric charge, dark matter (parameterized by PFDM parameter $\lambda_{\text{DM}}$), and nonlinear electrodynamics effects / QED effects (parameterized by nonlinear electrodynamics parameter $a$) correspondingly. In this figure, we present results on the trajectory of orbits and their precession angles for a selected angular momentum value $J=4M$. The variation trend of precession angles corresponding to other angular momentum values is similar to that for $J=4M$ presented in figure 5.